# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

### v1

@@ B.1. Basics
The first inequality is an application of the triangle inequality. The second inequality follows from Jensen’s inequality. Since $m\geq 1$, the function $x\mapsto x^{m}$ on $[0,\infty)$ is convex, and thus Jensen’s inequality implies
@@
Lemma 5 (Portmanteau lemma (Lemma 2.2 in Vaa (98))).
@@
Remark 1.
If $X_{n}$ converges to 00 in probability, then by Fact 2, we have
@@
B.2. Pseudo-Lipschitz Functions
Definition 2 (Pseudo-Lipschitz functions (BM, 11)).
@@ B.2. Pseudo-Lipschitz Functions
holds in general, Lemma 4 implies
@@
Hence, by Lemma 4, we have
@@
Appendix C Remaining proofs
C.1. Proof of Theorem 3
In this section we provide detailed proofs for the results sketched in Section 4, thereby completing the proof of Theorem 3.
Throughout, $\mathbb{E}[\cdot\mid X]$ denotes the conditional expectation with respect to the $\sigma$-algebra $\sigma(X)$. Since conditional expectations are only defined up to almost-sure equality, we omit “$\mathrm{a.s.}$” when writing “=$\stackrel{{\scriptstyle\mathrm{a.s.}}}{{=}}$” in this context.
C.1.1. Weak Convergence of the Dot Products
@@
Proposition 8.
Under the assumptions of Theorem 3, the vector $(p_{1},\dots,p_{r})$ converges in distribution to $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which is the Gaussian vector defined in Definition 1.
The proof of Proposition 8 relies on the next lemma, which is an application of Theorem 2 in BCRT (58).
Lemma 9.
@@
It is important to note that $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is always a deterministic value that is independ of $n$, and is fixed by the program architecture. According to the sampling rule explained in Section 2.1, the matrices $W^{i,j}$ and $W^{i^{\prime}j^{\prime}}$ are sampled independently whenever $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is zero. In particular, Assumption 1 gives
@@
Lemmas 1015 show that the sequence $\{X_{\alpha}\}_{\alpha\in[n]}$ satisfies the conditions of Lemma 9. The proof of Proposition 8 is completed by applying Lemma 9 to $S=\sum_{\alpha=1}^{n}X_{\alpha}/\sqrt{n}$ and then invoking the Cramér–Wold device.
Lemma 10.
@@
where the convergence follows from Lemma 16. Finally, Eq. (8) and Definition 1 imply that
@@
By a calculation similar to Lemma 12, we have
@@
×$\quad\times\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{1}}^{i_{2},j_{1}}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{2}}^{i_{2},j_{1}^{\prime}}\right)\left(\frac{1}{n}\sum_{\gamma_{3}=1}^{n}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{3}}^{i_{4},j_{2}}\right)\left(\frac{1}{n}\sum_{\gamma_{4}=1}^{n}x_{\gamma_{4}}^{i_{3},2}x_{\gamma_{4}}^{i_{4},j_{2}^{\prime}}\right)\right]$
@@
×$\quad\times\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j_{1}}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j_{1}^{\prime}}}\right]\mathbb{E}\left[Z^{x^{i_{3},1}}Z^{x^{i_{4},j_{2}}}\right]\mathbb{E}\left[Z^{x^{i_{3},2}}Z^{x^{i_{4},j_{2}^{\prime}}}\right],$
where the convergence follows from Lemma 16. Observe that this limit is equivalent to $\sigma_{*}^{4}$. ∎
Lemma 15.
@@
×$\quad\times\mathbb{E}\left[x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{2},1}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{4}}^{i_{4},1}x_{\gamma_{5}}^{i_{1},2}x_{\gamma_{6}}^{i_{2},2}x_{\gamma_{7}}^{i_{3},2}x_{\gamma_{8}}^{i_{4},2}\right].$
@@
Note that $C$ is bounded by the boundedness of $x^{i,j}$ for all $i\in[r]$ and $j\in[2]$ (see Definition 1). Then, we can write
@@
Thus, it remains only to prove that $\tilde{R}$ converges to 00 as $n\to\infty$. This can be achieved by showing that $\mathbb{E}(R_{\ell}^{4})$ converges to 00 for all $\ell\in[4]$. By Fact 1 and Lemma 7, for all $\ell\in[4]$, we know $R_{\ell}$ converges almost surely to 00. The continuous mapping theorem then yields
@@
To upgrade this to convergence in expectation, Facts 2 and 3 imply it suffices to show the existance of a constant $\delta>1$ that satisfies
@@
But since each $x^{i,j}$ and its infinite-width limit $Z^{x^{i,j}}$ are bounded, such a $\delta$ exists. Therefore, we conclude that $\mathbb{E}(R_{\ell}^{4})$ converges to 00 for all $\ell\in[4]$, and consequently, $\tilde{R}$ does as well. ∎
C.1.2. $S_{1}$ Converges to 0
@@
Fix a small $\epsilon>0$. Since $\mathring{\bm{p}}_{1:r}$ is a Gaussian vector (see Definition 1), it is tight. Hence there exists a compact set $K\subset\mathbb{R}^{r}$ that satisfies
@@
with some constant $C$. Furthermore, noting that $K$ is a Borel set, applying Lemma 5 and Proposition 8 gives
@@
Observe that $\tilde{\psi}_{\bm{a}_{1:r}}$ is bounded by the boundedness of $\psi$. We later show that $\tilde{\psi}_{\bm{a}_{1:r}}$ is also pseudo-Lipschitz (Lemma 17). Therefore, by Eq. (7) and Lemma 5, we have
@@
Moreover, noting that $f$ and $\psi$ are bounded and continuous (see Fact 4), we can apply the (uncoditional and conditional) bounded convergence theorem to show that $\Phi_{n}$ is continuous at every point. Therefore, we can apply Lemma 18 to show that $\sup_{\bm{a}_{1:r}\in K}\Phi_{n}(\bm{a}_{1:r})$ converges to 00 as $n\to\infty$.
Lemma 17.
@@ C.1.3. $S_{2}$ Converges to 0
where $Z^{\bm{g}^{1:M}}$ is independent of $\mathring{\bm{p}}_{1:r}$. By Lemma 5, this implies $S_{2}\to 0$.
C.2. Proof of Corollary 2
Let $\psi:\mathbb{R}^{J}\to\mathbb{R}$ be a bounded and Lipschitz function that satisfies $|\psi|\leq C$. Note that by Fact 4, it is also pseudo-Lipschitz. Define a bounded and continuous function $f:\mathbb{R}\to\mathbb{R}$ by
@@ C.2. Proof of Corollary 2
Then, Lemma 5 and Theorem 1 imply that
@@
holds as $n\to\infty$. Since the above convergence holds for all bounded and Lipschitz function $\psi$, by Lemma 5, this implies the desired convergence in distribution.

### v2

@@ C.1 Basics
The first inequality is an application of the triangle inequality. The second inequality follows from Jensen’s inequality. Since $m\geq 1$, the function $x\mapsto x^{m}$ on $0,\infty)$ is convex, and thus Jensen’s inequality implies
@@
Lemma C.2 (Portmanteau lemma (Lemma 2.2 in [29])).
@@
Remark C.1.
If $X_{n}$ converges to 0 in probability, then by Fact C.1, we have
@@
C.2 Pseudo-Lipschitz Functions
Definition C.1 (Pseudo-Lipschitz functions [6]).
@@ C.2 Pseudo-Lipschitz Functions
holds in general, Lemma C.1 implies
@@
Hence, by Lemma C.1, we have
@@
Appendix D Remaining proofs
D.1 Proof of Theorem 4.1
In this section we provide detailed proofs for the results sketched in Section 4, thereby completing the proof of Theorem 4.1.
Throughout, $\mathbb{E}[\cdot\mid X]$ denotes the conditional expectation with respect to the $\sigma$-algebra $\sigma(X)$. Since conditional expectations are only defined up to almost-sure equality, we omit $\mathrm{a.s.}$” when writing $\stackrel{{\scriptstyle\mathrm{a.s.}}}{{=}}$” in this context.
D.1.1 Weak Convergence of the Dot Products
@@
Proposition D.1.
Under the assumptions of Theorem 4.1, the vector $(p_{1},\dots,p_{r})$ converges in distribution to $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which is the Gaussian vector defined in Definition 3.1.
The proof of Proposition D.1 relies on the next lemma, which is an application of Theorem 2 in [5].
Lemma D.2.
@@
It is important to note that $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is always a deterministic value that is independent of $n$, and is fixed by the program architecture. According to the sampling rule explained in Section 2.1, the matrices $W^{i,j}$ and $W^{i^{\prime}j^{\prime}}$ are sampled independently whenever $d_{(i,j)}^{(i^{\prime},j^{\prime})}$ is zero. In particular, Assumption 3.1 gives
@@
Lemmas D.3D.8 show that the sequence $\{X_{\alpha}\}_{\alpha\in[n]}$ satisfies the conditions of Lemma D.2. The proof of Proposition D.1 is completed by applying Lemma D.2 to $S=\sum_{\alpha=1}^{n}X_{\alpha}/\sqrt{n}$ and then invoking the Cramér–Wold device.
Lemma D.3.
@@
where the convergence follows from Lemma D.9. Finally, Eq. (9) and Definition 3.1 imply that
@@
By a calculation similar to Lemma D.5, we have
@@
$\quad\times\mathbb{E}\left[\left(\frac{1}{n}\sum_{\gamma_{1}=1}^{n}x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{1}}^{i_{2},j_{1}}\right)\left(\frac{1}{n}\sum_{\gamma_{2}=1}^{n}x_{\gamma_{2}}^{i_{1},2}x_{\gamma_{2}}^{i_{2},j_{1}^{\prime}}\right)\left(\frac{1}{n}\sum_{\gamma_{3}=1}^{n}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{3}}^{i_{4},j_{2}}\right)\left(\frac{1}{n}\sum_{\gamma_{4}=1}^{n}x_{\gamma_{4}}^{i_{3},2}x_{\gamma_{4}}^{i_{4},j_{2}^{\prime}}\right)\right]$
@@
$\quad\times\mathbb{E}\left[Z^{x^{i_{1},1}}Z^{x^{i_{2},j_{1}}}\right]\mathbb{E}\left[Z^{x^{i_{1},2}}Z^{x^{i_{2},j_{1}^{\prime}}}\right]\mathbb{E}\left[Z^{x^{i_{3},1}}Z^{x^{i_{4},j_{2}}}\right]\mathbb{E}\left[Z^{x^{i_{3},2}}Z^{x^{i_{4},j_{2}^{\prime}}}\right],$
where the convergence follows from Lemma D.9. Observe that this limit is equivalent to $\sigma_{*}^{4}$. ∎
Lemma D.8.
@@
$\quad\times\mathbb{E}\left[x_{\gamma_{1}}^{i_{1},1}x_{\gamma_{2}}^{i_{2},1}x_{\gamma_{3}}^{i_{3},1}x_{\gamma_{4}}^{i_{4},1}x_{\gamma_{5}}^{i_{1},2}x_{\gamma_{6}}^{i_{2},2}x_{\gamma_{7}}^{i_{3},2}x_{\gamma_{8}}^{i_{4},2}\right].$
@@
Note that $C$ is bounded by the boundedness of $x^{i,j}$ for all $i\in[r]$ and $j\in[2]$ (see Definition 3.1). Then, we can write
@@
Thus, it remains only to prove that $\tilde{R}$ converges to 0 as $n\to\infty$. This can be achieved by showing that $\mathbb{E}(R_{\ell}^{4})$ converges to 0 for all $\ell\in[4]$. By Fact 3.1 and Lemma C.4, for all $\ell\in[4]$, we know $R_{\ell}$ converges almost surely to 0. The continuous mapping theorem then yields
@@
To upgrade this to convergence in expectation, Facts C.1 and C.2 imply it suffices to show the existance of a constant $\delta>1$ that satisfies
@@
But since each $x^{i,j}$ and its infinite-width limit $Z^{x^{i,j}}$ are bounded, such a $\delta$ exists. Therefore, we conclude that $\mathbb{E}(R_{\ell}^{4})$ converges to 0 for all $\ell\in[4]$, and consequently, $\tilde{R}$ does as well. ∎
D.1.2 $S_{1}$ Converges to 0
@@
Fix a small $\epsilon>0$. Since $\mathring{\bm{p}}_{1:r}$ is a Gaussian vector (see Definition 3.1), it is tight. Hence there exists a compact set $K\subset\mathbb{R}^{r}$ that satisfies
@@
with some constant $C$. Furthermore, noting that $K$ is a Borel set, applying Lemma C.2 and Proposition D.1 gives
@@
Observe that $\tilde{\psi}_{\bm{a}_{1:r}}$ is bounded by the boundedness of $\psi$. We later show that $\tilde{\psi}_{\bm{a}_{1:r}}$ is also pseudo-Lipschitz (Lemma D.10). Therefore, by Eq. (7) and Lemma C.2, we have
@@
Moreover, noting that $f$ and $\psi$ are bounded and continuous (see Fact C.3), we can apply the (uncoditional and conditional) bounded convergence theorem to show that $\Phi_{n}$ is continuous at every point. Therefore, we can apply Lemma D.11 to show that $\sup_{\bm{a}_{1:r}\in K}\Phi_{n}(\bm{a}_{1:r})$ converges to 0 as $n\to\infty$.
Lemma D.10.
@@ D.1.3 $S_{2}$ Converges to 0
where $Z^{\bm{g}^{1:M}}$ is independent of $\mathring{\bm{p}}_{1:r}$. By Lemma C.2, this implies $S_{2}\to 0$.
D.2 Proof of Corollary 3.2
Let $\psi:\mathbb{R}^{J}\to\mathbb{R}$ be a bounded and Lipschitz function that satisfies $|\psi|\leq C$. Note that by Fact C.3, it is also pseudo-Lipschitz. Define a bounded and continuous function $f:\mathbb{R}\to\mathbb{R}$ by
@@
Then, Lemma C.2 and Theorem 3.1 imply that
@@
holds as $n\to\infty$. Since the above convergence holds for all bounded and Lipschitz function $\psi$, by Lemma C.2, this implies the desired convergence in distribution.